A family currently live in an apartment whose monthly rent is $1050. They are thinking of buying a house which would cost $250,000. They plan to live in this house for 5 years and sell it at the end of the 5th year. They would put a downpayment of $25,000 and finance the balance through a mortgage at 3% interest rate. The mortgage is to be repaid in 5 annual installments (which include both principal and interest) at the end of each year for the next 5 years The house will have the following additional expenses: annual maintenance: $1500; Property taxes:$5500; Insurance: $1200. Assume they are in tax bracket of 20% and the price of home, rent and expenditure increases by 2% per year. Their opportunity cost or required rate of return is 5% per year. Note that property taxes are tax deductible and there no tax payable on capital gains. Use annual compounding for amortization schedule of mortgage.

Calculate the Post tax Mortgage Cost (principal repayment plus after tax interest cost) for year 4.

$44,078

$48,844

$48,566

$48,296
Answer #1
Yearly payment = [P × R × (1+R)^N ] / [(1+R)^N -1]
Using the formula:
Loan amount P $                                                          225,000
Rate of interest per period:
Annual rate of interest 3.000%
Frequency of payment = Once in 12 month period
Numer of payments in a year = 12/12 = 1
Rate of interest per period R 0.03 /1 = 3.0000%
Total number of payments:
Frequency of payment = Once in 12 month period
Number of years of loan repayment =                                                                          5
Total number of payments N 5 × 1 = 5
Period payment using the formula = [ 225000 × 0.03 × (1+0.03)^5] / [(1+0.03 ^5 -1]
Yearly payment = $                                                      49,129.78

Year 4 interest payment:

Period Beginning liability Uniform monthly payment Interest owed Principal payment Total owed at end of month
N A C B= A* 0.030000 D=C-B E=A-D
1            225,000.00     49,129.78             6,750.00     42,379.78      182,620.22
2            182,620.22     49,129.78             5,478.61     43,651.17      138,969.05
3            138,969.05     49,129.78             4,169.07     44,960.71        94,008.34
4              94,008.34     49,129.78             2,820.25     46,309.53        47,698.81

Interest paid in year 4 = 2,820.25

Total post tax cost = Principal paid + after tax interest

= 46,309.53 + 2,820.25 * (1-20% = $48,566

Answer is $$48,566